# Coherent Control of the Fine-Structure Qubit in a Single Alkaline-Earth Atom

### 19 Jan 2024 v1

@@ Abstract
We report on the first realization of a novel neutral atom qubit encoded in the metastable fine-structure states ${{}^{3}\rm{P}_{0}}$ and ${{}^{3}\rm{P}_{2}}$ of single ${}^{88}$r atoms trapped in an optical tweezer. Raman coupling of the qubit states promises rapid single-qubit rotations on par with the fast Rydberg-mediated two-body gates. We demonstrate preparation, read-out, and coherent control of the qubit. In addition to driving Rabi oscillations bridging an energy gap of more than 17 THz using a pair of phase-locked clock lasers, we also carry out Ramsey spectroscopy to extract the transverse qubit coherence time $T_{2}$. When the tweezer is tuned into magic trapping conditions, which is achieved in our setup by tuning the tensor polarizability of the ${{}^{3}\rm{P}_{2}}$ state via an external control magnetic field, we measure $T_{2}=1.2\,\mathrm{ms}$. A microscopic quantum mechanical model is used to simulate our experiments including dominant noise sources. We identify the main constraints limiting the observed coherence time and project improvements to our system in the immediate future. Our work opens the door for a so far unexplored qubit encoding concept for neutral atom based quantum computing.
@@
Here, we report on the first realization of a third possibility for qubit encoding in divalent atomic systems. Specifically, we exploit the two long-lived fine-structure states ${{}^{3}\rm{P}_{0}}$ and ${{}^{3}\rm{P}_{2}}$ (magnetic quantum number $m_{J}=0$) of the metastable triplet $5\rm{s}5\rm{p}$-manifold in the bosonic ${}^{88}$r atom [21]. The two qubit states are gapped by $17.419\,\mathrm{THz}$ and are coupled using a two-photon Raman transition via the intermediate $5\rm{s}6\rm{s}\,^{3}\rm{S}_{1}$ state (Fig. 1(a)). The fine-structure encoding comes with a number of notable benefits. First, the two-photon coupling promises single-qubit rotations on the 100 ns timescale, orders of magnitude faster than what has been demonstrated on the comparatively slow ${{}^{1}\rm{S}_{0}}$-${{}^{3}\rm{P}_{0}}$ optical clock qubit, bringing single-qubit gate times on par with the fast Rydberg-mediated two-qubit gates. Second, the small momentum transfer associated with the Raman drive substantially reduces recoil heating and provides clear advantages for coherent qubit transport. Third, the qubit features magic wavelength trapping near 540 nm, for which also Rydberg states are trapped via the polarizability of the ${}^{+}$ ionic core. Furthermore, exploiting the tunability of the tensor polarizability of the ${{}^{3}\rm{P}_{2}}$ state via an external magnetic field allows for finding a triple-magic trapping scenario, for which the two qubit states and the Rydberg level experience equal ac-Stark shifts, providing ideal situations for immediate next steps toward high-fidelity two-qubit entangling gates via single-photon coupling to Rydberg states [21, 22].
@@
At these conditions, the magic wavelength of the fine-structure qubit, which yields equal trap depth for ${{}^{3}\rm{P}_{0}}$ and ${{}^{3}\rm{P}_{2}}$ (intersection of the solid red and blue line in Fig. 1(c)), is about $4\,\mathrm{nm}$ shorter than our tweezer wavelength [24, 23]. For the tweezer power set during the Rabi oscillation measurement $P=1.45\,\mathrm{mW}$, we compute the differential light shift at the trap center $\delta U=U_{{{}^{3}}\rm{P}_{0}}-U_{{{}^{3}}\rm{P}_{2}}=-h\times 0.2\,\mathrm{MHz}$. Even though this yields a substantial non-magicness, we attribute the damping of the Rabi oscillations mainly to shot-to-shot fluctuations of the local Rabi frequency due to position fluctuations of the Raman beams relative to the tweezer. This is supported by a numerical simulation matched to the data (gray line). For this, we time-evolve a laser-driven two-level atom trapped in a 3D harmonic oscillator, taking into account level-dependent trap frequencies computed from the experimental tweezer parameters, finite temperature ($T=8\,\mathrm{\upmu K}$) and Monte Carlo sampling of the local Rabi frequency. The latter is taken from a Gaussian distribution with a standard deviation of $10\%$ around a mean value of $\Omega_{\rm{R}}=2\pi\times 84\,\mathrm{kHz}$.
@@
Figure 3: Magic wavelength tuning of the fine-structure qubit. Normalized Ramsey fringe contrast as a function of magnetic field angle $\phi$ for $|\vec{B}|=3\,\mathrm{G}$ (triangles) and $|\vec{B}|=8\,\mathrm{G}$ (circles) measured at fixed $t_{\rm{R}}=600\,\mathrm{\upmu s}$ and $t_{\rm{R}}=700\,\mathrm{\upmu s}$, respectively. For both datasets $P=46\,\mathrm{\upmu W}$. The solid line is a Gaussian to guide the eye. Insets show representative datasets for $\phi=9.5^{\circ}$ and $\phi=20^{\circ}$ taken at $|\vec{B}|=8\,\mathrm{G}$.
The dominant source of transverse qubit dephasing becomes evident when comparing Fig. 2(a) to a measurement recorded at a reduced tweezer power of $P=46\,\mathrm{\upmu W}$ shown in Fig. 2(b), for which we observe coherent oscillations at much longer times. This finding suggests that dephasing is dominated by non-magic qubit trapping in combination with finite temperature. Again our numerical simulations (solid lines in Fig. 2(a) and (b)), for which we compute the quantum mechanical time evolution as before, but now incorporating the non-standard Ramsey protocol, strongly support this claim. More specifically, the results in Fig. 2(a) and Fig. 1(d) assume the same temperature and Rabi frequency fluctuations, while the simulation data in Fig. 2(b) use $T=1.4\,\mathrm{\upmu K}$ as expected for adiabatic ramping to the shallow tweezer setting after trap loading. In summary, the analysis shows that while Rabi oscillations are most sensitive to Raman beam fluctuations, the Ramsey measurements are largely independent of that, and provide an excellent measure for trap-induced qubit dephasing.
Figure 4: Coherence time of the fine-structure qubit. (a) Ramsey signal normalized to the inital fitted fringe constrast at magic angle $\phi=9.5^{\circ}$ for $|\vec{B}|=8\,\mathrm{G}$ measured in the shallow tweezer ($P=46\,\mathrm{\upmu W}$). The gray lines are sinusoidal fits to the data to extract the temporal decay of the contrast. (b) Contrast extracted from fits to data as shown in (a) for $\phi=0^{\circ}$ in the deep tweezer (squares), and for $\phi=0^{\circ}$ (triangles) and $\phi=9.5^{\circ}$ (crosses,circles) in the shallow trap at $|\vec{B}|$ as indicated in the legend. Dashed lines show fits based on a Gaussian envelope to extract the qubit coherence time $T_{2}$.
@@
A Ramsey measurement recorded at this magic trapping angle is shown in Fig. 4(a) for $|\vec{B}|=8\,\mathrm{G}$. Noteably, tuning $\phi$ allows us to clearly observe qubit coherence for times exceeding 1.5 ms, providing an improvement of more than an order of magnitude compared to the situation presented in Fig 2(a). For a more quantitative analysis, we fit the Ramsey data with sinusoidal functions (solid lines), from which we extract the fringe contrast as a function of $t_{\rm{R}}$. The result of that procedure is presented in Fig. 4(b) (circles), and is compared to scenarios at $\phi=0^{\circ}$ for both the shallow and the deep tweezer setting (triangles and squares).
@@
The influence of the longitudinal field components on $\delta U(x,y)$ yields a quadratic increase with the distance from the trap center along the direction perpendicular to the input polarization vector, before it becomes maximal near the trap waist with a value $\delta U_{\rm{max}}\approx h\times 1.7\,\mathrm{kHz}$. We may estimate the effect on qubit dephasing by averaging $\delta U(x,y)$ over a thermal Gaussian distribution in the focal plane with the temperature obtained previously ($T=1.4\,\mathrm{\upmu K}$), and find a corresponding timescale of a few milliseconds, in fair agreement with our observed $T_{2}$ time.
@@
In conclusion, we have demonstrated a novel fine-structure qubit for neutral atom quantum computing with divalent atoms. We have studied the coherence properties of this qubit in an optical tweezer trap and measured a coherence time of up to 1.2 ms when the trap is tuned into magic trapping conditions, exploiting the tensor polarizability of the ${{}^{3}\rm{P}_{2}}$ state. We found that the observed $T_{2}$ time is limited by the polarization landscape in the tweezer focus. Improved in-trap cooling of the qubit, ideally into the three-dimensional ground state, should significantly increase the coherence time toward the tens of milliseconds range. This has to be compared to the exceptionally fast single-qubit rotations, which can be readily achieved via the reported Raman coupling. Specifically, increasing the Raman intensity into the milliwatt regime and using a tighter beam waist enables single-qubit rotations on the 100 ns timescale. This provides exciting prospects for achieving $\sim 10^{5}$ gate operations within the qubit coherence time, which is comparable with the so far studied clock or nuclear spin qubit implementations in divalent atoms.
@@
We thank M. S. Safronova for providing polarizability data, and the QRydDemo consortium and the Quantum Länd team for fruitful discussions. We acknowledge funding from the Federal Ministry of Education and Research (BMBF) under the grants QRydDemo, CiRQus, MUNIQC-Atoms, and the Horizon Europe programme HORIZON-CL4-2021-DIGITAL-EMERGING-01-30 via the project 101070144 (EuRyQa).

### 13 Mar 2024 v2

@@ Abstract
We report on the first realization of a novel neutral atom qubit encoded in the spin-orbit coupled metastable states ${{}^{3}\rm{P}_{0}}$ and ${{}^{3}\rm{P}_{2}}$ of a single ${}^{88}$r atom trapped in an optical tweezer. Raman coupling of the qubit states promises rapid single-qubit rotations on par with the fast Rydberg-mediated two-body gates. We demonstrate preparation, read-out, and coherent control of the qubit. In addition to driving Rabi oscillations bridging an energy gap of more than 17 THz using a pair of phase-locked clock lasers, we also carry out Ramsey spectroscopy to extract the transverse qubit coherence time $T_{2}$. When the tweezer is tuned into magic trapping conditions, which is achieved in our setup by tuning the tensor polarizability of the ${{}^{3}\rm{P}_{2}}$ state via an external control magnetic field, we measure $T_{2}=1.2\,\mathrm{ms}$. A microscopic quantum mechanical model is used to simulate our experiments including dominant noise sources. We identify the main constraints limiting the observed coherence time and project improvements to our system in the immediate future. Our work opens the door for a so far unexplored qubit encoding concept for neutral atom based quantum computing.
@@
Here, we report on the first realization of a third possibility for qubit encoding in divalent atomic systems. Specifically, we exploit the two long-lived fine-structure states ${{}^{3}\rm{P}_{0}}$ and ${{}^{3}\rm{P}_{2}}$ (magnetic quantum number $m_{J}=0$) of the metastable triplet $5\rm{s}5\rm{p}$-manifold in the bosonic ${}^{88}$r atom [21]. The two qubit states are gapped by $17.419\,\mathrm{THz}$ and are coupled using a two-photon Raman transition via the intermediate $5\rm{s}6\rm{s}\,^{3}\rm{S}_{1}$ state (Fig. 1(a)). The fine-structure encoding comes with a number of notable benefits. First, the two-photon coupling promises single-qubit rotations on the 100 ns timescale, orders of magnitude faster than what has been demonstrated on the comparatively slow ${{}^{1}\rm{S}_{0}}$-${{}^{3}\rm{P}_{0}}$ optical clock qubit, bringing single-qubit gate times on par with the fast Rydberg-mediated two-qubit gates. Compared to sub-microsecond gates realized in alkali atoms via Raman coupling utilizing vector light shifts [22], our qubit yields favorable heating properties for very large single-photon detunings, similar to recent explorations of lanthanides for spin-orbit coupled Fermi gases [23]. Second, the small momentum transfer associated with the Raman drive substantially reduces recoil heating and provides clear advantages for coherent qubit transport. Third, the qubit features magic wavelength trapping near 540 nm, for which also Rydberg states are trapped via the polarizability of the ${}^{+}$ ionic core. Furthermore, exploiting the tunability of the tensor polarizability of the ${{}^{3}\rm{P}_{2}}$ state via an external magnetic field allows for finding a triple-magic trapping scenario, for which the two qubit states and the Rydberg level experience equal ac-Stark shifts, providing ideal situations for immediate next steps toward high-fidelity two-qubit entangling gates via single-photon coupling to Rydberg states [21, 24]. Finally, we note that spin-orbit encoding for qubit implementation is also at the heart of various other platforms such as quantum dots [25], color centers [26], or even large molecular complexes [27].
@@
At these conditions, the magic wavelength of the fine-structure qubit, which yields equal trap depth for ${{}^{3}\rm{P}_{0}}$ and ${{}^{3}\rm{P}_{2}}$ (intersection of the solid red and blue line in Fig. 1(c)), is about $4\,\mathrm{nm}$ shorter than our tweezer wavelength [29, 28]. For the tweezer power set during the Rabi oscillation measurement $P=1.45\,\mathrm{mW}$, we compute the differential light shift at the trap center $\delta U=U_{{{}^{3}}\rm{P}_{0}}-U_{{{}^{3}}\rm{P}_{2}}=-h\times 0.2\,\mathrm{MHz}$. Even though this yields a substantial non-magicness, we attribute the damping of the Rabi oscillations mainly to shot-to-shot fluctuations of the local Rabi frequency due to position fluctuations of the Raman beams relative to the tweezer. This is supported by a numerical simulation matched to the data (gray line). For this, we time-evolve a laser-driven two-level atom trapped in a 3D harmonic oscillator, taking into account level-dependent trap frequencies computed from the experimental tweezer parameters, finite temperature ($T=8\,\mathrm{\upmu K}$) and Monte Carlo sampling of the local Rabi frequency. The latter is taken from a Gaussian distribution with a standard deviation of $10\%$ around a mean value of $\Omega_{\rm{R}}=2\pi\times 84\,\mathrm{kHz}$. This corresponds to large beam position fluctuations (1.4 $\upmu$m) since we focus the tweezer and Raman beams via independent objectives in the current constrained setup, which can be significantly improved by focusing both via the same objective, enabling fluctuations on the 100 nm level and hundreds of Rabi cycles.
@@
Figure 3: Magic wavelength tuning of the fine-structure qubit. Normalized Ramsey fringe contrast as a function of magnetic field angle $\phi$ for $|\vec{B}|=3\,\mathrm{G}$ (triangles) and $|\vec{B}|=8\,\mathrm{G}$ (circles) measured at fixed $t_{\rm{R}}=600\,\mathrm{\upmu s}$ and $t_{\rm{R}}=700\,\mathrm{\upmu s}$, respectively. For both datasets $P=46\,\mathrm{\upmu W}$. The solid line is a Gaussian to guide the eye. Insets show representative datasets for $\phi=9.5^{\circ}$ and $\phi=20^{\circ}$ taken at $|\vec{B}|=8\,\mathrm{G}$. The contrast is extracted from the amplitude of a sinusoidal fit to the Ramsey oscillations and normalized to the maximum value obtained at the magic angle.
The dominant source of transverse qubit dephasing becomes evident when comparing Fig. 2(a) to a measurement recorded at a reduced tweezer power of $P=46\,\mathrm{\upmu W}$ shown in Fig. 2(b), for which we observe coherent oscillations at much longer times. This finding suggests that dephasing is dominated by non-magic qubit trapping in combination with finite temperature [30]. Again our numerical simulations (solid lines in Fig. 2(a) and (b)), for which we compute the quantum mechanical time evolution as before, but now incorporating the non-standard Ramsey protocol, strongly support this claim. More specifically, the results in Fig. 2(a) and Fig. 1(d) assume the same temperature and Rabi frequency fluctuations, while the simulation data in Fig. 2(b) use $T=1.4\,\mathrm{\upmu K}$ as expected for adiabatic ramping to the shallow tweezer setting after trap loading. In summary, the analysis shows that while Rabi oscillations are most sensitive to Raman beam fluctuations, the Ramsey measurements are largely independent of that, and provide an excellent measure for trap-induced qubit dephasing.
Figure 4: Coherence time of the fine-structure qubit. (a) Ramsey signal normalized to the inital fitted fringe constrast at magic angle $\phi=9.5^{\circ}$ for $|\vec{B}|=8\,\mathrm{G}$ measured in the shallow tweezer ($P=46\,\mathrm{\upmu W}$). The gray lines are sinusoidal fits to the data to extract the temporal decay of the contrast. (b) Contrast extracted from fits to data as shown in (a) for $|\vec{B}|=3\,\mathrm{G},\phi=0^{\circ}$ in the deep tweezer (squares) and for the shallow tweezer at $|\vec{B}|=3\,\mathrm{G},\phi=0^{\circ}$ (triangles), $|\vec{B}|=3\,\mathrm{G},\phi=9.5^{\circ}$ (crosses), and $|\vec{B}|=8\,\mathrm{G},\phi=9.5^{\circ}$ (circles) with $T_{2}$ values of 36(2) $\upmu$s, 203(17)$\upmu$s, 676(70) $\upmu$s and 1236(82) $\upmu$s, respectively. Dashed lines show fits based on a Gaussian envelope to extract the qubit coherence time $T_{2}$.
@@
A Ramsey measurement recorded at this magic trapping angle is shown in Fig. 4(a) for $|\vec{B}|=8\,\mathrm{G}$. Notably, tuning $\phi$ allows us to clearly observe qubit coherence for times exceeding 1.5 ms, providing an improvement of more than an order of magnitude compared to the situation presented in Fig 2(a). For a more quantitative analysis, we fit the Ramsey data with sinusoidal functions (solid lines), from which we extract the fringe contrast as a function of $t_{\rm{R}}$. The result of that procedure is presented in Fig. 4(b) (circles), and is compared to scenarios at $\phi=0^{\circ}$ for both the shallow and the deep tweezer setting (triangles and squares).
@@
The influence of the longitudinal field components on $\delta U(x,y)$ yields a quadratic increase with the distance from the trap center along the direction perpendicular to the input polarization vector, before it becomes maximal near the trap waist with a value $\delta U_{\rm{max}}\approx h\times 1.7\,\mathrm{kHz}$. We may estimate the effect on qubit dephasing by averaging $\delta U(x,y)$ over a thermal Gaussian distribution in the focal plane with the temperature obtained previously ($T=1.4\,\mathrm{\upmu K}$), and find a corresponding timescale of a few milliseconds, in fair agreement with our observed $T_{2}$ time. We note the possibility to rephase this trap-induced decoherence via spin-echo, for which we expect to reach irreversible coherence times in the range of tens of milliseconds, dictated by heating during the spin-echo protocol [30].
@@
In conclusion, we have demonstrated a novel fine-structure qubit for neutral atom quantum computing with divalent atoms. We have studied the coherence properties of this qubit in an optical tweezer trap and measured a coherence time of up to 1.2 ms when the trap is tuned into magic trapping conditions, exploiting the tensor polarizability of the ${{}^{3}\rm{P}_{2}}$ state. We found that the observed $T_{2}$ time is limited by the polarization landscape in the tweezer focus. Improved in-trap cooling of the qubit, ideally into the three-dimensional ground state, should significantly increase the coherence time toward the tens of milliseconds range. Additionally, we propose an elegant route to largely eliminate the differential trap depth variations. One may choose a tweezer wavelength for which the magic angle $\phi=90\,\mathrm{{}^{\circ}}$. Considering the focal spot as a superposition of plane waves, it is apparent that the field vectors are all orthogonal to the magnetic field axis, up to the much smaller transverse component. Notably, one finds such conditions in the strontium atom for a wavelength of approximately 755 nm, for which we predict a reduction of the maximum differential trap depth as shown in Fig 5(b) by two orders of magnitude.
Our target coherence times of tens of milliseconds and longer have to be compared to the exceptionally fast single-qubit rotations, which can be readily achieved via the reported Raman coupling. Specifically, increasing the Raman intensity into the milliwatt regime and using a tighter beam waist enables single-qubit rotations on the 100 ns timescale. This provides exciting prospects for achieving $\sim 10^{5}$ gate operations within the qubit coherence time, which is comparable with the so far studied clock or nuclear spin qubit implementations in divalent atoms.
Finally, high-fidelity readout of the qubit can be implemented in the future via coherent transfer starting from ${{}^{3}\rm{P}_{0}}$, either directly to ${{}^{1}\rm{S}_{0}}$ via a three-photon scheme, or to ${{}^{3}\rm{P}_{1}}$ with two photons, followed by spontaneous decay. Moreover, transfer into a tweezer or lattice system at the clock-magic wavelength at about 813 nm [17] would even provide simultaneous magic trapping of the clock- and the fine-structure qubit, enabling coherent state transfer between two qubits encoded in the same atom. In that context the clock transition may serve as an ideal storage with interesting prospects for mid-circuit read-out and feedback in a qudit setting.
@@
In a study performed in parallel to ours, similar results have been achieved with atoms in an optical lattice [34]. We thank M. S. Safronova for providing polarizability data, and the QRydDemo consortium and the Quantum Länd team for fruitful discussions. We acknowledge funding from the Federal Ministry of Education and Research (BMBF) under the grants QRydDemo, CiRQus, MUNIQC-Atoms, and the Horizon Europe programme HORIZON-CL4-2021-DIGITAL-EMERGING-01-30 via the project 101070144 (EuRyQa).